The sum of Roger's marks in Science and math is 156. His mark in math exceeds his mark in science by 10.
If x represents Roger's mark in science, which of the following equations could be used to determine Roger's marks in science and math? PLZ HELP? and give me a detailed explanations, I struggle with these kind of questions alot

Answer by stanbon(75887)

The sum of Roger's marks in Science and math is 156. His mark in math exceeds his mark in science by 10.
If x represents Roger's mark in science, which of the following equations could be used to determine Roger's marks in science and math?
------
Scores Equation::: s + m = 156
m = s + 10
------
------
Substitute to get:
s + s + 10 = 156
2s + 10 = 156
=====================
Cheers,
Stan H.
=====================